OpenStudy (anonymous):

How to solve 0.3(x+3) = 0.10x + 1.80

OpenStudy (anonymous):

0.3(x+3) = 0.3 x + 0.9 = 0.10x + 1.80 can you do the rest?

OpenStudy (anonymous):

um im a little confused with what u did

OpenStudy (anonymous):

What is 0.3 * 3 =??

OpenStudy (anonymous):

0.9

OpenStudy (anonymous):

a(b+c) = ab+ac

OpenStudy (anonymous):

This is known as distributive law.

OpenStudy (anonymous):

oh you just distributed sorry yeah i get it

OpenStudy (anonymous):

now what do i do?

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

\[ 0.3 x + 0.9 = 0.10x + 1.80 \implies 0.7x =- 0.9 \]

OpenStudy (anonymous):

Combining the like terms.

OpenStudy (anonymous):

subtract 0.3 x from both sides

OpenStudy (anonymous):

okay now you just divide

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

Ok Good work :D @FoolForMath

OpenStudy (anonymous):

wait but i typed this into a website and it gave me 4.5 it wouldn't show the steps though and when i do it like this it doesn't come out 4.5

OpenStudy (anonymous):

LOL. Multiply the numerator and denominator by 10 and then try

OpenStudy (anonymous):

0.3(x+3) = (0.10x + 1.80) do the parentheses make a difference now??

OpenStudy (mertsj):

Take your original equation and multiply each term by 10 to remove all decimal points and then use the distributive property: \[10\times.3(x+3)=10\times.10x+10\times1.80\] \[3(x+3)=1x+18\] \[3x+9=x+18\] \[2x=9\] \[x=\frac{9}{2}\]

OpenStudy (anonymous):

why would you multiply by 10 though???

OpenStudy (anonymous):

To get rid of decimal

OpenStudy (mertsj):

makes it easier to eliminate decimal points

OpenStudy (anonymous):

Basically If a = b+c Then 10a = 10b+10c

OpenStudy (anonymous):

true dat

OpenStudy (anonymous):

so now the answer is 4.5

OpenStudy (anonymous):

Wait, I was being a FOOL, my last step was incorrect, \[ 0.3 x + 0.9 = 0.10x + 1.80 \implies 0.2x = 0.9 \implies x= \frac 9 2 \]

OpenStudy (anonymous):

okay lol thanks you guys!

OpenStudy (anonymous):

Glad to help!